# Competing effect of disorder on phase separation in active systems: from facilitated to hindered growth and rough interfaces

### v1

@@ Abstract
We investigate the impact of random pinned disorder on a collection of self-propelled particles. To achieve this, we construct a continuum model by formulating the coupled hydrodynamic equations for slow variables: local density and momentum density of particles. The disorder in the system acts as pinning sites, effectively immobilizing the particles that come into contact with them. Our numerical results reveal that weak disorder leads to phase separation in the system at density and activity lower than the typical values for motility induced phase separation. We construct a phase diagram using numerical simulations as well as linearized approximation in the plane of activity and packing fraction of particles at weak disorder densities. As disorder densities rise in the system, kinetic processes slow down, while at high disorder densities, the system becomes heterogeneous and eventually undergoes kinetic arrest. The structure factor tail deviates from Porod’s law, indicating increased roughness at domain interfaces under strong disorder. Furthermore, we analyze the fractal dimension of the interface as a function of disorder density, highlighting the increasing irregularity of phase-separated domains.
@@ I Introduction
Nonequilibrium living and lab designed active systems for example, cells, tissues, living organisms and autonomous robots Ramaswamy (2010, 2017); Marchetti et al. (2013); Vrugt and Wittkowski (2024); Volpe et al. (2022) composed of many self-propelled agents unveil intriguing collective behavior across a wide range of length and time scales. The study of active matter systems became an interesting area of research for many years due to the emergent behaviors such as pattern formation Denk and Frey (2020); Hallatschek et al. (2023); Jena and Mishra (2024), nonequilibrium disorder to order transitions Kumar et al. (2024), anomalous fluctuations Jena and Mishra (2023); Ramaswamy et al. (2003), and interesting behaviour in different medium and confinement etc. within these systems, that are not present in corresponding analogous equilibrium systems. Another captivating characteristic of such systems is Motility Induced Phase Separation (MIPS) that resembles the passive liquid-gas phase separation but occurs in absence of any attractive interactions Fily and Marchetti (2012); Cates and Tailleur (2015) at much lower packing densities.
@@
Majority of studies of active matter in theory and experiments are focused on the systems in homogeneous or clean environment Jülicher et al. (2018); Sanchez et al. (2012); Zöttl and Stark (2016); Alston et al. (2022); Aranson (2022); Gachelin et al. (2014); Pattanayak and Mishra (2018); Sampat and Mishra (2021); Takatori and Brady (2016); Fily and Marchetti (2012); Wittkowski et al. (2014); Stenhammar et al. (2013). But, in natural active matter systems, inhomogeneity or disorder is present intrinsically. Various types of disorder or inhomogeneities are observed in natural systems that arise from multiple factors. This can be present in the form of spatial (geometric variation), temporal (moving obstacles), chemical composition and, biological inhomogeneity (bio-diversity) etc. Exploring and understanding such systems with disorder is essential to realize the complexity of natural systems and the systems’ response to internal as well as external stimuli.
There has been extensive studies on active systems in clean environment, while the exploration of systems with disorder is relatively less. However, there exist a few prior studies on systems possessing disorder Kumar and Mishra (2020); Das et al. (2018); Chen et al. (2022); Vahabli and Vicsek (2023); Pinçe et al. (2016); Ro et al. (2021); Goswami and Chakrabarti (2022); Mishra and Mishra (2023). These studies explore the effects of both quenched and annealed disorder in active systems, highlighting their impact on the dynamics of active particles by examining transport properties and diffusivity. In certain cases, disorder enhances collective behavior in active polar particles, while in active scalar particles, it has been reported to induce long-range order Ro et al. (2021). The findings Kumar et al. (2021); Sándor et al. (2017); Reichhardt and Reichhardt (2018) reveal that the disorder can lead to the emergence of distinct phases. In most studies, disorder acts as either obstacles placed physically in the space or a random field disorder Das et al. (2018); Chepizhko and Peruani (2013); Ro et al. (2021); Morin et al. (2017). Apart from some studies, very little is still known about the phase behavior of the active systems in the presence of disorder which acts like pinning sites and immobilize the particles. This kind of fixed obstacles can be noticed in many natural system i.e, motion of bacteria in porous media Drescher et al. (2011), migrating cells encountering collagen fibers Wolf and Friedl (2011), collective cell migration through tissue environment Petrie et al. (2009) etc.
@@
Our primary focus in the present work is on the effect of pinned disorder on the phase separation and kinetics of collection of self-propelled particles, setting it apart from previous works. The study will be beneficial for observing the dynamical behavior of the organism in the presence of disorder in natural system.
@@
In the present work, we introduce disorder, as pinning sites in a system consists of a collection of active particles, by constructing a continuum model by formulating the coupled hydrodynamic equations for slow variables: local density and momentum density of particles. The pinning sites are modeled in such a manner that they effectively make the self-propelled speed of the active particles zero. We explore the phase diagram to gain insight into the phenomenon of phase separation in the presence of disorder. We observed phase separation at density and activity lower than the critical values for a clean system Fily and Marchetti (2012); Redner et al. (2013), a result further confirmed by linearized hydrodynamic calculations. Although the pinned disorder eases the phase separation, the kinetics of phase separation slows down at lower disorder and becomes arrested at high disorder densities. The structure factor tail exhibits a deviation from Porod’s law, suggesting enhanced roughness at domain interfaces in the presence of strong disorder. Additionally, our analysis of the fractal dimension of the interface as a function of disorder density reveals irregularity in phase-separated domains.
@@ II Model and Numerical Methodology
We develop a continuum model to study a collection of self-propelled particles on a two dimensional substrate with the presence of random pinning sites. We formulate a minimal hydrodynamic coupled equations for conserved local density field $\rho(\bm{r},t)$ and local orientation order parameter $\bm{p}(\bm{r},t)$ and further we define the momentum density $\bm{P}(\bm{r},t)=\rho(\bm{r},t)\bm{p}(\bm{r},t)$. The hydrodynamic equation for the density $\rho(\bm{r},t)$ is:
@@
here, $v(\rho)=v_{0}(1-\lambda\rho)$ denoted as the effective velocity of each particle, is influenced by the local density of the particles and the presence of pinned sites in the system. Its dependence of local density is adopted from previous observations that the local clustering suppresses the motile nature of particles Fily and Marchetti (2012). Further, we introduce the pinned obstacles or disorder in such a manner there are few randomly selected regions in the substrate such that $v(\rho)=0$ at those points. The number density of such points on the substrate is the density of pinning sites in the system is defined as $\rho_{d}$.
Eq.1 is a continuity equation where, the first term on the right side of the equation accounts for the active current generated by self-propulsion and the second term represents the diffusion current with diffusion constant $D_{\rho}$. Eq. 2 is similar to the equation introduced by Fily et.al Fily and Marchetti (2012) in the context of self-propelled particles neglecting the nonlinearites in $\bm{P}$. The first and second terms on the right hand side of Eq. 2 represents the polarization decays at rate $\nu_{r}$ and is convected by pressure-like gradients $\sim\bm{\nabla}(v(\rho)\rho)$. The third term represents the diffusion in orientation. And $\bm{f}_{\rho}$ and $\bm{f}_{\bm{P}}$ are the Gaussian white noise in density and polarization equation respectively having strengths $\Delta_{\rho,\bm{P}}$;
@@
where, $\alpha\equiv(\rho,\bm{P})$ and $(i,j)$ can take two values $x$ and $y$ representing the two cartesian coordinates. The intrinsic time scale $\tau$ and intrinsic length scale $l_{0}$ are defined as $\nu_{r}^{-1}$ and $\sqrt{D_{\rho}/\nu_{r}}$ respectively. The Eq. 1, 2 are rescaled by the $\tau$ and $l_{0}$. Numerical integration of Eq. 1 and 2 is performed with homogeneous initial density with mean $\rho_{0}$ and random ${\bm{P}}$. We vary the mean density from $\rho_{0}$ = 0.2 to 1.0 and self-propulsion speed from $v_{0}=0$ to $7$ in a box of size $K\times K$ with periodic boundary condition in the both directions. To investigate the impact of disorder in the system, we vary the disorder density $\rho_{d}$ from 00 to $0.5$. The integration is performed using Euler’s scheme Bally and Talay (1996) with $\Delta x=1.0l_{0}$ and $\Delta t=0.1\tau$. In the system under consideration, the parameters are fixed as follows: $\nu_{r}=k=D_{\rho}=1.0$ and $\lambda=0.9$. For these choice of parameters the initial homogeneous state becomes unstable and system phase separates for $\rho_{0}>1/(2\lambda)$ as reported in Fily and Marchetti (2012). We performed numerical simulations for simulation time, $t=2\times 10^{5}$ and for system sizes $K=256-1024$ and averaging is performed over $50$ independent realizations.
Figure 1: The panels (a-d) showcase the local fluctuation $\delta\rho$ at across four panels , each representing system with different disorder densities $\rho_{d}=0.0,0.05,0.1,0.5$ respectively at $t=200$. Within each panel, multiple figures are plotted from left to right for self-propulsion speed $v_{0}=0.2,1,3,5,7$ keeping the mean density $\rho_{0}=0.6$ fixed. The color on the heatmap represents the value of $\delta\rho$. The results are obtained for systems having $K=256$.
@@
Figure 3: The plot(a) presents semi-log $x-$ plot of $\Delta\rho$ vs. $\rho_{d}$ with error bars. The inset shows a linear plot of $\Delta\rho$ vs. $\rho_{d}$ depicting the non-monotonicity with respect to disorder. The plot(b) illustrates the phase diagram in the $\rho_{0}v_{0}^{2}-v_{0}^{2}$ plane for different disorder densities: $\rho_{d}=0.0,0.001$ respectively. The blue (circles) and red (squares) solid lines are the boundary drawn from the numerical simulation for $\rho_{d}=0.0,0.001$ which is also mentioned in the legend. Additionally, the green solid line and magenta dotted line represent the analytical boundary obtained from linearized calculation. The shaded light-orange region shows the extra regime of phase separation in the presence of disorder. The plot (c) showcases $\Delta\rho$ vs. $v_{0}$ depicting the transition from non-phase separation to phase separation for $\rho_{d}=0.0,0.001,0.01$.
Figure 4: The panels (a-d) showcase the snapshots of time series of the local density field $\rho$ at t = 5, 50, 150 and 200 across the columns in sequence. Each panel from top to bottom, depicts the snapshots for system with $\rho_{d}=0,0.05,0.3,0.5$, respectively. The color in the heatmap represent the magnitude of local density at each lattice point. The results are obtained from simulating a $256\times 256$ system.
III Result
@@
To investigate the effect of disorder on the system, we begin by plotting snapshots of normalized local density fluctuations (${\delta\rho}$) defined as ${\delta\rho}({\bm{r}},t)=\frac{\rho({\bm{r}},t)-\rho_{0}}{\rho_{0}}$. We examine the result for a range of ($v_{0}$, $\rho_{0}$) values in systems with varying density of disorder $\rho_{d}$. Figs. 1(a-d) illustrate snapshots of local densities fluctuations across four panels, showcasing the effects of varying $\rho_{d}$ and fixing $\rho_{0}=0.6$. Panel (a) represents a system without disorder ($\rho_{d}=0$) for different values of $v_{0}=0.2,1,3,5,7$. Panels (b), (c), and (d) display the corresponding results for systems with disorder densities of $\rho_{d}$ = 0.05, 0.1, and 0.5, respectively. Fig. 1(a) shows that in a clean system, phase separation begins at $v_{0}>3$ with the formation of small domains, indicating the onset of Motility-Induced Phase Separation $(MIPS)$, consistent with previous studies on active Brownian particles Fily and Marchetti (2012). For the systems with finite density of disorder, ${\delta\rho}$ starts to develop finite values at $v_{0}=3$ as can be seen in Fig. 1(b) for $\rho_{d}=0.05$. Thus, the density inhomogeneity develops at a speed below that for the clean system. As the numbers of pinning sites increases, as shown in Fig 1(c) for $\rho_{d}=0.1$, the density inhomogeneity starts to appear at lower $v_{0}$. But, after a certain $\rho_{d}$ the system does not phase separate and instead shows the heterogeneous density as shown in Fig. 1(d) for $\rho_{d}=0.5$. Similarly in Fig. 2(a-d), we show the fluctuations in local density for fixed $v_{0}=6$ , varying mean particle densities $\rho_{0}$ , for different disorder densities.
@@
The density of disorder is the same as for the Fig. 1(a-d). This allows us to examine how both $\rho_{0}$ and disorder influence the system’s behavior. We observed that the system undergoes phase separation at ${\rho_{0}}\geq 0.6$ for the clean system ($\rho_{d}=0$) in Fig. 2(a). As suggested by previous studies for clean systems Fily and Marchetti (2012), $MIPS$ occurs once a critical density is reached. However, as the disorder in the system increases, phase separation occurs even at lower values of $\rho_{0}$ as shown in Fig.2(b-c). Therefore, the phase diagram in the ${\rho_{0}}-\rho_{d}$ plane demonstrates that disorder promotes density inhomogeneity in the system for weak disorder. Similar to Fig.1(d), higher disorder makes system heterogeneous again as shown in Fig.2(d).
@@
Based on the observation from the snapshots shown in Figs. 1 and 2, it can be concluded that the system with finite $\rho_{d}$ shows the phase separating domains at self propulsion speed $v_{0}$ and mean density $\rho_{0}$ smaller than that for the clean system $\rho_{d}=0$. Hence the presence of small number of pinning sites drives the phase separation at lower activity and mean density. Further, the system with high density of disorder shows the formation of heterogeneous structures suppressing phase separation.
@@
To analyze it thoroughly, in Fig. 3(a) we show the plot of density phase separation order parameter $\Delta\rho$ (PSOP) vs. $\rho_{d}$, where $\Delta\rho=\frac{1}{K^{2}}<{\sum}_{\bm{r}}|\delta\rho(\bm{r},t)|>$, where <..><..>< . . > mean average over time in the steady state and over different realizations. $N$ is the number of lattice points in the system, $\Delta\rho$ reflects the amount of phase separation in the system. We find $\Delta\rho$ shows the non-monotonic feature as we increase the disorder density. We fix the self-propelled speed $v_{0}$ and $\rho_{0}$ to values 4.0 and 0.6 respectively, such that the homogeneous state is stable so that $\Delta\rho=0$ for the clean system. As we introduce disorder, $\Delta\rho$ shows a jump to finite value and again decreases for the large disorder densities as shown in Fig. 3(a)(inset). A very small disorder is enough to make the homogeneous state unstable as shown in Fig. 3(a)(main panel) (on semi-$\log x$- scale).
@@
We now investigate the mechanisms responsible for disorder-induced phase separation at low disorder densities and the transition to a heterogeneous phase at high disorder densities. In the low disordered environment, the accumulation around the obstacles could be a result of particles being immobilized upon reaching close proximity, which leads to effectively zero velocity causing the particles to gather in that location. Additionally, when new particles approach, the existing ones act as obstacles and contributing to the reduction of their velocity. As a result of this effect, clustering around the obstacles and domains formation is observed. Since, the phase separation is driven by the disorder in the system, we refer to it as Disorder-Induced Phase Separation ($DIPS$). However, for high $\rho_{d}$, the particles get immobilized at most of the sites. There is no scope for movement of the particles and they get stuck immediately around their nearest pinning sites. As the particles immobilize immediately, they can not accumulate in sufficient amount to form large clusters. Although the diffusion in density term is non-zero, but it will simply diffuse the particles from one place to another and pinned sites will act like cold regions with small accumulation of density around that sites. As a result, heterogeneous structures are formed in the system.

### v2

@@ Abstract
We investigate the impact of random pinned disorder on a collection of self-propelled particles. To achieve this, we construct a continuum model by formulating the coupled hydrodynamic equations for slow variables: local density and momentum density of particles. The disorder in the system acts as pinning sites, effectively immobilizing the particles that come into contact with them. Our numerical results reveal that weak disorder leads to phase separation in the system at density and activity lower than the typical values for motility induced phase separation. We construct a phase diagram using numerical simulations as well as linearized approximation in the plane of activity and packing fraction of particles at weak disorder densities. On increasing disorder density the system shows the Micro phase separation, while at large disorder densities, the system becomes heterogeneous and eventually undergoes kinetic arrest. The structure factor tail deviates from the Porod’s law, indicating increased roughness at domain interfaces under strong disorder. Furthermore, we analyze the fractal dimension of the interface as a function of disorder density, highlighting the increasing irregularity of phase-separated domains. We also found that disorder significantly suppresses number fluctuations in the system.
preprint: APS/123-QED
@@ I Introduction
Nonequilibrium living and lab designed active systems for example, cells, tissues, living organisms and autonomous robots [1, 2, 3, 4, 5] are composed of many self-propelled agents unveil intriguing collective behavior across a wide range of length and time scales. The study of active matter systems became an interesting area of research for many years due to the emergent behaviors such as pattern formation [6, 7, 8], nonequilibrium disorder to order transitions [9], anomalous fluctuations [10, 11], and interesting behavior in different medium and confinement etc., that are not present in corresponding equilibrium systems. Another captivating characteristic of such systems is Motility Induced Phase Separation (MIPS) that resembles the passive liquid-gas phase separation but occurs in absence of any attractive interactions [12, 13], [14, 15] at much lower packing densities.
@@
The majority of studies of active matter in theory and experiments are focused on the systems in homogeneous or clean environments [16, 17, 18, 19, 20, 21, 22, 23, 24, 12, 25, 26, 27]. But, in natural active matter systems, inhomogeneity or disorder is present intrinsically. Various types of disorder or inhomogeneities are observed in natural systems that arise from multiple factors. This can be present in the form of spatial (geometric variation), temporal (moving obstacles), chemical composition and, biological inhomogeneity (bio-diversity) etc. Exploring and understanding such systems with disorder is essential to realize the complexity of natural systems and the systems’ response to internal as well as external stimuli.
There have been extensive studies on active systems in clean environments, while the exploration of systems with disorder is relatively less. However, there exist a few prior studies on systems possessing disorder [28, 29, 30, 31, 32, 33, 34, 35, 36]. These studies explore the effects of both quenched and annealed disorder in active systems, highlighting their impact on the dynamics of active particles by examining transport properties and diffusivity. In certain cases, disorder enhances collective behavior in active polar particles, while in active scalar particles, it has been reported to induce long-range order [33]. The works by C. Reichhardt and C. J. O. Reichhardt have made significant contributions to the study of systems in disordered media over many years. Using microscopic methods, their work explores a variety of scenarios, including systems with periodic arrays of obstacles and inhomogeneous substrates [37, 38, 39, 40]. Their findings [37, 38] demonstrate that disorder can give rise to distinct dynamical phases. Notably, they observe phenomena such as pinning, ratchet effects, clogging, collective behavior, and pattern formation [41, 42, 43] etc.
In most of the other studies, disorder acts as either obstacles placed physically in the space or a random field disorder [29, 44, 33, 45, 46]. Apart from some studies, very little is still known about the phase behavior of the active systems in the presence of disorder which acts like pinning sites and immobilize the particles. The pinning sites as a disorder can be more clearly understood from a biological point of view. Pinned or “frozen” sites correspond to heterogeneous environments where cells or active agents cannot move and get stuck due to the mechanical trapping. There are plenty of examples of such biological systems that naturally contain such immobile, obstructive patches. This kind of fixed obstacles can be noticed in many natural system i.e, motion of bacteria in porous media [47], migrating cells encountering collagen fibers [48], collective cell migration through tissue environment [49] etc.
@@
Our primary focus in the present work is on the effect of spatial heterogeneity, modeled as pinned disorder, on phase behavior, phase separation kinetics, and resulting morphology of active matter systems using a field theoretic approach, which sets it apart from previous works. The study will be beneficial for observing the dynamical behavior of the organism in the presence of disorder in natural systems.
@@
In the present work, we introduce disorder, as pinning sites in a system consisting of a collection of active particles, by constructing a continuum model by formulating the coupled hydrodynamic equations for slow variables: local density and momentum density of particles. The pinning sites are modeled in such a manner that they effectively make the self-propelled speed of the active particles zero. We explore the phase diagram to gain insight into the phenomenon of phase separation in the presence of disorder. We observed phase separation at density and activity lower than the critical values for a clean system [12, 50], a result further confirmed by linearized hydrodynamic calculations. Although, the pinned disorder eases phase separation, the phase separation weakens with increasing disorder density. At moderate disorder densities Micro phase separation is found whereas the arrested kinetics is found at high disorder densities. The structure factor tail exhibits a deviation from the Porod’s law, suggesting enhanced roughness at domain interfaces in the presence of strong disorder. Additionally, our analysis of the fractal dimension of the interface as a function of disorder density reveals irregularity in phase-separated domains.
@@ II Model and Numerical Methodology
We develop a continuum model to study a collection of self-propelled particles on a two dimensional substrate with the presence of random pinning sites. We formulate a minimal hydrodynamic coupled equations for conserved local density field $\rho(\bm{r},t)$ and local orientation order parameter $\bm{p}(\bm{r},t)$ and further, we define the momentum density $\bm{P}(\bm{r},t)=\rho(\bm{r},t)\bm{p}(\bm{r},t)$. The hydrodynamic equation for the density $\rho(\bm{r},t)$ is:
@@
Here, $v(\rho)=v_{0}(1-\lambda\rho)$, effective velocity of each particle, is influenced by the local density of the particles and the presence of pinned sites in the system. A pinned disorder site corresponds to a rectangular region of size $\Delta x\times\Delta x$ and these regions are chosen randomly initially and then their spatial arrangement is quenched with time. The velocity dependence on the local density is adopted from previous observations that the local clustering suppresses the motile nature of particles [12]. Further, we introduce the pinned obstacles or disorder in such a manner that there are few randomly selected regions in the substrate such that $v(\rho)=0$ at those points, still the particle can diffuse because the noise $\bm{f}_{\rho}$ is not zero. Here, we mean the particle can only diffuse and lose its motility and thermal fluctuations can still lead them to diffuse. The number density of such points on the substrate is the density of pinning sites in the system, and is defined as $\rho_{d}$.
Eq.1 is a continuity equation where, the first term on the right side of the equation accounts for the active current generated by self-propulsion and the second term represents the diffusion current with diffusion constant $D_{\rho}$. Eq. 2 is similar to the equation introduced by Fily et.al [12] in the context of self-propelled particles neglecting the nonlinearites in $\bm{P}$. The first and second terms on the right hand side of Eq. 2 represent the polarization decay at rate $\nu_{r}$ and it’s convection by pressure-like gradients $\sim\bm{\nabla}(v(\rho)\rho)$ respectively. The third term represents diffusion in orientation. And $\bm{f}_{\rho}$ and $\bm{f}_{\bm{P}}$ are the Gaussian white noise in density and polarization equations respectively having strengths $\Delta_{\rho,\bm{P}}$;
@@
where, $\alpha\equiv(\rho,\bm{P})$ and $(i,j)$ can take two values $x$ and $y$ representing the two cartesian coordinates. The noise we have taken $\bm{f}_{\rho}$ and $\bm{f}_{\bm{P}}$ are additive in nature. For simplification in the model, we take the noise as additive which gives random kick and does not depend on the local density. The intrinsic time scale $\tau$ and intrinsic length scale $l_{0}$ are defined as $\nu_{r}^{-1}$ and $\sqrt{D_{\rho}/\nu_{r}}$ respectively. Eq. 1, 2 are rescaled by the $\tau$ and $l_{0}$. Numerical integration of Eq. 1 and 2 is performed with homogeneous initial density with mean $\rho_{0}$ and random $\bm{P}$}. We vary the mean density from $\rho_{0}$ = 0.2 to 0.9 and self-propulsion speed from $v_{0}=0$ to $7$ in a box of size $K\times K$ with periodic boundary condition in the both directions. To investigate the impact of disorder in the system, we vary the disorder density $\rho_{d}$ from 0 to $0.5$. The integration is performed using Euler’s scheme [51] with $\Delta x=1.0l_{0}$ and $\Delta t=0.1\tau$. In the system under consideration, the parameters are fixed as follows: $\nu_{r}=k=D_{\rho}=1.0$ and $\lambda=0.9$. For these choice of parameters, the initial homogeneous state becomes unstable and system phase separates for $\rho_{0}>1/(2\lambda)$ as reported in [12]. We performed numerical simulations for simulation time, $t=2\times 10^{5}$ and for system sizes K=2561024K=256-1024. Further, the averaging is performed over $50$ independent realizations to get better statistics.
Figure 1: The panels (a-d) showcase the local fluctuation $\delta\rho$ at across four panels , each representing system with different disorder densities $\rho_{d}=0.0,0.05,0.1,0.5$ respectively at $t=200$. Within each panel, multiple figures are plotted from left to right for self-propulsion speed $v_{0}=0.2,1,3,5,7$ keeping the mean density $\rho_{0}=0.7$ fixed. The color on the heatmap represents the value of $\delta\rho$. The results are obtained for systems having $K=256$.
@@
Figure 3: The plot(a) presents semi-log $x$- plot of $\Delta\rho$ vs. $\rho_{d}$ with error bars. The inset shows a linear plot of $\Delta\rho$ vs. $\rho_{d}$ depicting the non-monotonicity with respect to disorder. The plot(b) illustrates the phase diagram in the $\rho_{0}v_{0}^{2}-v_{0}^{2}$ plane for different disorder densities: $\rho_{d}=0.0,0.001$ respectively. The blue (circles) and red (squares) solid lines are the boundary drawn from the numerical simulation for $\rho_{d}=0.0,0.001$ which is also mentioned in the legend. Additionally, the green solid line and magenta dotted line represent the analytical boundary obtained from linearized calculation. The shaded light-orange region shows the extra regime of phase separation in the presence of disorder. The plot (c) showcases $\Delta\rho$ vs. $v_{0}$ depicting the transition from non-phase separation to phase separation for $\rho_{d}=0.0-0.2$.
III Results
@@
To investigate the effect of disorder on the system, we begin by plotting snapshots of normalized local density fluctuations ($\delta\rho$}) defined as ${\delta\rho}({\bm{r}},t)=\frac{\rho({\bm{r}},t)-\rho_{0}}{\rho_{0}}$. We examine the result for a range of ($v_{0}$, $\rho_{0}$) values in systems with varying density of disorder $\rho_{d}$. Figs. 1(a-d) illustrate snapshots of local density fluctuations across four panels, showcasing the effect of varying $\rho_{d}$ and fixing $\rho_{0}=0.7$. These snapshots are taken at fixed time $t=200$ (simulation time $t=2000$) which does not correspond to a steady state, it is in a transient state. Panel (a) represents a system without disorder ($\rho_{d}=0$) for different values of $v_{0}=0.2,1,3,5,7$. Panels (b), (c), and (d) display the corresponding snapshots for systems with disorder densities of $\rho_{d}$ = 0.05, 0.1, and 0.5, respectively. Fig. 1 Panel (a) shows that in a clean system, phase separation begins at $v_{0}>3$ with the formation of small domains, indicating the onset of Motility-Induced Phase Separation $(MIPS)$, consistent with previous studies on active Brownian particles [12]. For the systems with non-zero disorder density, the local density fluctuations $\delta\rho$} shows non-zero values at $v_{0}=3$ as can be seen in Fig. 1 Panel (b) for $\rho_{d}=0.05$, indicating phase separation in the system. Thus, the density inhomogeneity or phase separation develops at a speed below that for the clean system. As the numbers of pinning sites increases, as shown in Fig 1 Panel (c) for $\rho_{d}=0.1$, the density inhomogeneity starts to appear at lower $v_{0}$. But, after a certain $\rho_{d}$ the system does not phase separate and instead shows the heterogeneous density as shown in Fig. 1 Panel (d) for $\rho_{d}=0.5$.
@@
Similarly, in Fig. 2 Panel (a-d), we show the fluctuations in local density for fixed $v_{0}=6$ , varying mean particle densities $\rho_{0}$ , for different disorder densities. The density of disorder is the same as for the Fig. 1 Panel (a-d). This allows us to examine how both $\rho_{0}$ and disorder influence the system’s behavior. We observed that the system undergoes phase separation at ${\rho_{0}}>0.6$ for the clean system ($\rho_{d}=0$) in Fig. 2 Panel (a). As suggested by previous studies for clean systems [12], $MIPS$ occurs once a critical density is reached. However, as the disorder in the system increases, phase separation occurs even at lower values of $\rho_{0}$ as shown in Fig. 2 Panel (b-c). Therefore, the diagram demonstrates that disorder promotes density inhomogeneity in the system for weak disorder. Similar to Fig. 1 Panel (d), higher disorder makes system heterogeneous again as shown in Fig.2 Panel(d).
@@
Based on the observation from the snapshots shown in Figs. 1 and 2, it can be concluded that systems with finite $\rho_{d}$ shows the phase separating domains at self propulsion speed $v_{0}$ and mean density $\rho_{0}$ smaller than that for the clean system $\rho_{d}=0$. Hence, the presence of small number of pinning sites drives the phase separation at lower activity and mean density. Further, the system with high density of disorder shows the formation of heterogeneous structures suppressing phase separation.
@@
To analyze it thoroughly, in Fig. 3(a) we show the plot of density phase separation order parameter $\Delta\rho$ $(PSOP)$ vs. $\rho_{d}$, where $\Delta\rho=\frac{1}{K^{2}}<{\sum}_{\bm{r}}|\delta\rho(\bm{r},t)|>$, where <..><..> mean average over time in the steady state and over different realizations. We calculated the $PSOP$ with respect to time and we define the steady state when the $PSOP$ does not significantly change with respect to time. Here, $\Delta\rho$ reflects the amount of phase separation in the system. We fix the self-propelled speed $v_{0}$ and $\rho_{0}$ to values 4.2 and 0.7 respectively, such that the homogeneous state is stable so that $\Delta\rho=0$ for the clean system. As we introduce disorder, $\Delta\rho$ shows a jump to finite value and again decreases for the large disorder densities monotonically as shown in Fig. 3(a)(inset). A very small disorder is enough to make the homogeneous state unstable as shown in Fig. 3(a)(main panel) (on semi-$\log x$- scale).
@@
We now investigate the mechanism responsible for disorder-induced phase separation at low disorder densities and the transition to a heterogeneous phase at high disorder densities. In the low disorder environment, the accumulation around the obstacles could be a result of particles being immobilized upon reaching close proximity, which leads to effectively zero velocity causing the particles to gather at that location. Additionally, when new particles approach, the existing ones act as obstacles and contributing to the reduction of their velocity. As a result of this effect, clustering around the obstacles and domains formation is observed. Since, the phase separation is driven by the disorder in the system, we refer to it as Disorder-Induced Phase Separation ($DIPS$). The movie MV1 (MV1) shows how the pinning sites (represented as black dots) act as nucleation sites and accelerate the formation of cluster of the active particles around it. The color bar in the movie depicts the local density field $\rho(r)$ of the active particles. To further quantify it, we have also calculated the correlation between active particles and disorder sites $C_{\rho\rho_{d}}(t)$. The mathematical definition of $C_{\rho\rho_{d}}$ is given in Appendix A It is a measure of likelihood of finding a cluster of particle near a disorder site. The positive $C_{\rho\rho_{d}}$ means clustering of particle close to disorder sites, or low density of particle, where there is no disorder. The plot is generated for low disorder density $\rho_{d}=0.05$ as shown in Fig A1. From the plot, the correlation is seen to start with a small positive value, then gradually increase, and eventually saturate at late times. This behavior indicates that, initially, clusters nucleate in the vicinity of the disorder sites. After sufficient time, the saturation of the correlation suggests that $MIPS$ becomes the dominant mechanism. We discussed this in details while explaining Fig. 4 (a). However, for high $\rho_{d}$, the particles get immobilized at most of the sites. There is no scope for movement of the particles and they get stuck immediately around their nearest pinning sites. As the particles immobilize immediately, they can not accumulate in sufficient amount to form large clusters. Although, the diffusion in density term is non-zero, but it simply diffuses the particles from one place to another and pinned sites act like cold regions with small accumulation of density around that sites. As a result, heterogeneous structures are formed in the system.
Phase Diagram: